# Few-Shot Recognition via Stage-wise Retrieval-Augmented Finetuning

### v2

@@ E Analysis of Data Augmentation Methods
F Further Analyses on SWAT
We compare different classifier initialization methods for SWAT (Table 12). Results show that initializing with text embedding yields better performance than random initialization. In addition. we show the detailed performance of classifier retraining for each dataset in Table 13. The rare classes of the Aircraft dataset show significant performance gains (>10%) after classifier retraining, demonstrating the efficacy of classifier retraining with few-shot data in mitigating domain gaps and imbalanced distribution. In addition, we include the detailed ablation of SWAT components on each dataset in Table 14. Results show that applying CutMix [74] and classifier retraining effectively mitigate the domain gap and imbalanced distribution problem, verifying the design of SWAT.
@@ F Further Analyses on SWAT
Table 12: Comparison of classifier initialization methods in SWAT. We compare the final test accuracy by initializing the classifier before stage-1 end-to-end finetuning in different ways. Initializing classifier weights with text embedding features leads to better performance than random initialization. [32] explains that using randomly initialized classifier weights to finetune the model can distort the features of pretrained model, leading to worse finetuning performance. Throughout this work, we use prompts in [42] to initialize classifier weights in SWAT.
@@
Table 14: Ablation study on important components in our SWAT. We show the detailed performance improvements by each component for each dataset in our SWAT. Finetuning on simply combined retrieved and few-shot data underperfroms finetuning solely on few-shot data (8-shot and 16-shot, with or without CutMix), due to the large domain gap and imbalanced distribution in retrieved data. However, further applying CutMix and classifier retraining improves the test accuracy significantly, confirming their effectiveness in mitigating the domain gap and imbalanced distributions. We also compare the performance of few-shot finetuning with and without CutMix data augmentation. The results indicate more few-shot data yields more improvements, likely due to stronger data augmentation.
@@
Table 15: Impact of retrieval size (number of images per class) on the performance of SWAT. We show the performance of SWAT on each dataset using different numbers of retrieved images. We highlight the best number in bold and underline the second best. Importantly, we find that retrieving 10 images per class works the best for Flowers, EuroSAT, DTD, and Cars datasets. This is probably because LAION-400M contains limited images that matches these downstream concepts and simply retrieving more will include more noisy images and more imbalanced distributions, which hurt the training performance. We list the performance of previous state-of-the-art few-shot recognition method CLAP [57] in the table for comparison with our SWAT.
@@
G Analysis of Retrieved Data
We show the imbalanced distribution of retrieved data for all nine datasets in Fig. 7. We report the total number of retrieved images per dataset with increasing retrieval size (images per class) in Table 16. With increasing retrieval size, the total number of retrieved images increases less significantly due to the limited presentation of many downstream concepts in the pretraining datasets (e.g. LAION [53, 54]). To address this issue, we suggest future work to retrieve relevant images from diverse data sources, e.g. other datasets or the Internet [33].
Additionally, we compare SWAT’s performance on different retrieval size in Table 15. Results show that SWAT saturates at 500 images per class for 4-shot and 8-shot cases, and at 300 for 16-shot. Notably, for Flowers, EuroSAT, DTD, and Cars, retrieving 10 images per class yields the best results, likely due to improved data balance and the removal of noisy images (Fig. 10). Future work can study how to enhance the balance and quality of retrieved data. Fig. 8 shows more examples of retrieved images for each dataset.
@@ G Analysis of Retrieved Data
Figure 9: Visual comparison between downstream DTD images and the retrieved images (from LAION-400M [53]) for various DTD concepts. Clearly, large domain gap exists between the two data resources regarding styles, backgrounds, semantics, etc. In addition, the retrieved images only have a partial region depicting the texture, contrasting to the few-shot image which are “almost entirely filled with a texture” according to DTD’s strict data collection rules [11]. We suggest future work to explore better retrieval methods that are closely aligned with downstream data distribution, e.g., by referring to the data collection/annotation rules provided in the data annotation guidelines of a downstream task.
Figure 10: Examples of noisy retrieved images (from LAION-400M [53]) filtered by T2I thresholding. We show that string-matching-based retrieval (by searching image captions that contain any concept synonyms) can retrieve noisy images that could compromise the learning of downstream concepts, e.g., the bird eggs or the distribution map of bird species (first row). Using text-to-image (T2I) filtering helps remove such noisy images, and improve the performance of SWAT (Table 10). We choose a T2I threshold of 0.25 for our experiment, similar to that used in the curation of LAION [53, 54]. We highlight the T2I cosine similarity under each image and concept synonyms in the each image caption.
Table 16: Total number of retrieved images for each dataset under different retrieval size. With a larger retrieval size (number of retrieved images per class), we observe a diminished increase in the total number of retrieved images. This is because many downstream concepts have limited presence in the pretraining set (LAION-400M [53, 54]). See the imbalanced distribution of each dataset in Fig. 7.
imgs per class Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet
@@
H Analysis of Classifier Retraining
In Fig. 11, we show the final test accuracy after retraining the classifier across varying epoch numbers. For all datasets, accuracy remains stable with additional epochs. The small standard deviations across three runs with different random seeds confirm that stage-2 classifier retraining with few-shot data does not suffer from overfitting.
Refer to caption Refer to caption
Refer to caption Refer to caption
@@ I Code and Instructions
Requirements. Running our code requires some common packages. We installed Python and most packages through Anaconda. A few other packages might not be installed automatically, such as clip, open_clip_torch, img2dataset, torchvision, and PyTorch, which are required to run our code. We provide detailed instructions for building environment in file ENV.md. Below are the versions of Python and PyTorch used in our work.
@@
    Create conda environment and install required packages.

### v3

@@ E Analysis of Data Augmentation Methods
Impact of Mixing Ratio. We further explore the impact of the mixing ratio between retrieved and few-shot data within a batch when applying CutMix augmentation. Results in Fig. 7 shows that SWAT achieves the best performance when applying a “natural ratio” by combining retrieved data and few-shot annotated data without sophisticated resampling methods. This encourages future work to explore better mixed sample data augmentation methods.
Figure 7: Comparison of final accuracy with varying few-shot ratio in a batch. Our SWAT adopts a “natural ratio” by combining retrieved data and few-shot annotated data without sophisticated resampling methods. The natural ratio is 3%, meaning 3% data in each batch is from the few-shot data. Results show that the “natural ratio” (red dashed line) performs better than either increasing the ratio (which reduces data diversity) or decreasing it (which increases domain gap).
F Validating the Design of SWAT
Table 12: Comparison of ImageNet accuracy and training time cost of different stage-2 training strategies. We experiment by finetuning the stage-1 trained model on 16-shot data from ImageNet following different training strategies. Results show that finetuning only the classifier (as done in SWAT) improves the rare class accuracy significantly more than finetuning the visual encoder only. In addition, the training time cost of retraining the classifier is much less than finetuning the visual encoder. Moreover, finetuning both the visual encoder and classifier achieves further improvement over SWAT, likely due to the insufficient representation learning in stage 1 with only 50 training epochs. We denote this scenario as SWAT+ and report its performance across all datasets in Fig. 5, Table 7 and Table 8.
FT encoder FT classifier Avg common rare time
acc after stage-1 67.1 68.3 56.1
67.6+0.5 68.3+0.0 61.2+5.1 0.5 mins
67.4+0.3 68.5+0.2 57.3+1.2 15 mins
69.3+2.2 70.1+1.8 62.0+5.9 15 mins
Table 13: Comparison of different finetuning methods. We compare SWAT with state-of-the-art probing-based and finetuning-based methods using the same training data (a mix of retrieved and few-shot data). We experiment with the T2I-filtered retrieved data for each dataset (cf. Table 10). We use the same set of hyperparameters in Section B for all methods except using a larger batch size of 256 for FLYP following [17]. Results show that finetuning-based methods largely outperform probing-based methods, indicating the necessity of finetuning the visual encoder to learn better representation. In addition, ensembling the finetuned model with the zero-shot model (WiSE-FT with $\alpha=0.5$ [69]) leads to much worse accuracy than standard finetuning, likely because the zero-shot OpenCLIP model struggles to recognize these fine-grained concepts [52]. Finally, SWAT outperforms other finetuning methods, validating its effectiveness in mitigating domain gaps and imbalanced distribution issues in retrieved data. We highlight the best number in bold and underline the second best.
method Semi-Aves Flowers Aircraft EuroSAT DTD mean accuracy
(shots) 4 8 16 4 8 16 4 8 16 4 8 16 4 8 16 4 8 16
linear probing [49] ICML’24 49.8 52.4 54.4 86.9 89.4 92.8 34.6 35.8 38.2 68.0 78.2 82.4 61.7 65.5 68.9 60.2 64.3 67.3
CMLP [36] CVPR’23 49.2 51.9 53.6 87.0 89.3 92.9 34.1 35.4 37.8 70.1 79.4 83.5 61.3 64.8 68.6 60.3 64.2 67.3
REAL-Linear [43] CVPR’24 51.0 52.5 54.3 85.0 86.4 88.7 31.2 31.8 33.8 66.5 73.4 76.2 62.2 64.7 67.4 59.2 61.8 64.1
standard FT [49] ICML’24 55.2 57.6 60.4 89.4 92.8 95.5 48.9 51.2 53.0 83.3 88.3 92.8 61.5 65.6 70.3 67.7 71.1 74.4
WiSE-FT [69] CVPR’22 51.7 53.2 56.1 82.1 84.6 87.0 32.2 33.2 34.0 77.4 85.2 87.4 64.1 66.7 69.4 61.5 64.6 66.8
FLYP [17] CVPR’23 56.0 57.7 59.6 88.1 91.1 94.4 47.9 49.7 51.2 75.4 83.3 90.6 63.1 67.4 70.3 66.1 69.2 72.6
SWAT (ours) 58.6 61.3 63.8 91.0 94.7 97.5 55.5 58.1 62.6 84.6 89.2 93.7 63.0 67.6 71.6 70.5 74.2 77.8
Ablation of Stage-2 Training Strategy. To validate the design of stage-2 classifier retraining in SWAT, we compare the performance of different stage-2 training strategies in Table 12. Results show that retraining only the classifier achieves significantly larger accuracy improvement on rare classes than retraining only the visual encoder, validating its effectiveness in mitigating imbalanced distribution. Furthermore, we find that retraining both the visual encoder and classifier improves further over SWAT by 1$\sim$2%. We hypothesize that this is due to the insufficient representation learning in stage 1 with only 50 training epochs (recall that we follow realistic evaluation protocol that do not use validation set to tune hyperparameters). A supporting evidence is found in Fig. 8 where we show that longer training in stage 1 generally yields better final accuracy. We denote this strategy as SWAT+ and report its performance across all datasets in Fig. 5, Table 7 and Table 8. Considering the comparable performance and much less training time cost, we adopt classifier retraining for stage 2 in our SWAT.
Comparison with SOTA Finetuning Methods. Table 13 shows that our SWAT outperforms recent probing-based or finetuning-based methods using the same retrieved and few-shot data. SWAT also outperforms recent ensembling-based [69] and contrastive finetuning [17] methods, highlighting the effectiveness of our proposed stage-wise training in mitigating domain gap and imbalanced distribution issues.
Figure 8: Comparison of final accuracy with increasing stage-1 training epochs. Results show that increasing stage-1 training epochs generally increases final accuracy slightly, without overfitting issues. This is likely due to the improved representation learning. We set stage-1 training epochs to 50 for all datasets, following the realistic FSR setup that does not tune hyperparameters using a large validation set.
G Further Analyses on SWAT
Impact of Stage-1 Training Epochs. We compare the final accuracy with a varying number of epochs for stage-1 end-to-end finetuning in Fig. 8. Results show that longer training generally yields better performance due to improved representation learning. Please note that our realistic FSR setup does not allow using a validation set to tune training epochs. Our paper sets the number of training epochs to 50 for all datasets (cf. Section B).
@@ G Further Analyses on SWAT
Table 14: Comparison of classifier initialization methods in SWAT. We compare the final test accuracy by initializing the classifier before stage-1 end-to-end finetuning in different ways. Initializing classifier weights with text embedding features leads to better performance than random initialization. [33] explains that using randomly initialized classifier weights to finetune the model can distort the features of pretrained model, leading to worse finetuning performance. Throughout this work, we use prompts in [43] to initialize classifier weights in SWAT. Subscripts mark the performance improvement compared with random initialization.
@@
Classifier Initialization. We compare different classifier initialization methods for SWAT (Table 14). Results show that initializing with text embedding yields better performance than random initialization.
Retraining Classifier does not Overfit. In Fig. 9, we show the final test accuracy after retraining the classifier across varying epoch numbers. For all datasets, accuracy remains stable with increasing epochs. The small standard deviations across three runs with different random seeds confirm that stage-2 classifier retraining with few-shot data does not suffer from overfitting.
Refer to caption Refer to caption
Refer to caption Refer to caption
More Detailed Experimental Results. In addition. we show the detailed performance of classifier retraining for each dataset in Table 15. The rare classes of the Aircraft dataset show significant performance gains (>10%) after classifier retraining, demonstrating the efficacy of classifier retraining with few-shot data in mitigating domain gaps and imbalanced distribution. In addition, we include the detailed ablation of SWAT components on each dataset in Table 16. Results show that applying CutMix [76] and classifier retraining effectively mitigate the domain gap and imbalanced distribution problem, verifying the design of SWAT.
@@
Table 16: Ablation study on important components in our SWAT. We show the detailed performance improvements by each component for each dataset in our SWAT. Finetuning on simply combined retrieved and few-shot data underperforms finetuning solely on few-shot data (8-shot and 16-shot, with or without CutMix), due to the large domain gap and imbalanced distribution in retrieved data. However, further applying CutMix and classifier retraining improves the test accuracy significantly, confirming their effectiveness in mitigating the domain gap and imbalanced distributions. We also compare the performance of few-shot finetuning with and without CutMix data augmentation. The results indicate more few-shot data yields more improvements, likely due to stronger data augmentation.
@@
Table 17: Impact of retrieval size (number of images per class) on the performance of SWAT. We show the performance of SWAT on each dataset using different numbers of retrieved images. We highlight the best number in bold and underline the second best. Importantly, we find that retrieving 10 images per class works best for Flowers, EuroSAT, DTD, and Cars datasets. This is probably because LAION-400M contains limited images that match these downstream concepts and simply retrieving more will include more noisy images and more imbalanced distributions, which hurt the training performance. We list the performance of the previous state-of-the-art few-shot recognition method CLAP [59] in the table for comparison with our SWAT.
@@
H Analysis of Retrieved Data
Imbalances of Retrieved Data. We show the imbalanced distribution of retrieved data for all nine datasets in Fig. 10. We report the total number of retrieved images per dataset with increasing retrieval size (images per class) in Table 18. With increasing retrieval size, the total number of retrieved images increases less significantly due to the limited presentation of many downstream concepts in the pretraining datasets (e.g. LAION [55, 56]). To address this issue, we suggest future work to retrieve relevant images from diverse data sources, e.g. other datasets or the Internet [34]. Fig. 11 shows more examples of retrieved images for each dataset.
Impact of Retrieval Sizes. Additionally, we compare SWAT’s performance on different retrieval sizes in Table 17. Results show that SWAT saturates at 500 images per class for 4-shot and 8-shot cases and at 300 for 16-shot. Notably, for Flowers, EuroSAT, DTD, and Cars, retrieving only 10 images per class yields the best results, likely due to improved data balance and the exclusion of noisy images (Fig. 13). Future work can study how to enhance the balance and quality of retrieved data.
@@ H Analysis of Retrieved Data
Figure 12: Visual comparison between downstream DTD images and the retrieved images (from LAION-400M [55]) for various DTD concepts. Clearly, a large domain gap exists between the two data resources regarding styles, backgrounds, semantics, etc. In addition, the retrieved images only have a partial region depicting the texture, contrasting to the few-shot images which are “almost entirely filled with a texture” according to DTD’s strict data collection rules [11]. We suggest future work to explore better retrieval methods that are closely aligned with downstream data distribution, e.g., by referring to the data collection/annotation rules provided in the data annotation guidelines of a downstream task.
Figure 13: Examples of noisy retrieved images (from LAION-400M [55]) filtered by T2I thresholding. We show that string-matching-based retrieval (by searching image captions that contain any concept synonyms) can retrieve noisy images that could compromise the learning of downstream concepts, e.g., the bird eggs or the distribution map of bird species (first row). Using text-to-image (T2I) filtering helps remove such noisy images and improve the performance of SWAT (Table 10). We choose a T2I threshold of 0.25 for our experiment, similar to that used in the curation of LAION [55, 56]. We highlight the T2I cosine similarity and concept synonyms for each image.
Table 18: Total number of retrieved images for each dataset under different retrieval sizes. With a larger retrieval size (number of retrieved images per class), we observe a diminished increase in the total number of retrieved images. This is because many downstream concepts have limited presence in the pretraining set (LAION-400M [55, 56]). See the imbalanced distribution of each dataset in Fig. 10.
images / class Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet
@@ I Code and Instructions
Requirements. Running our code requires some common packages. We installed Python and most packages through Anaconda. A few other packages might not be installed automatically, such as clip, open_clip_torch, img2dataset, torchvision, and PyTorch, which are required to run our code. We provide detailed instructions for building the environment in file ENV.md. Below are the versions of Python and PyTorch used in our work.
@@
    Create a conda environment and install the required packages.